$SPAD/src/input robidoux.input

Nicolas Robidoux and Timothy Daly

December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool robidoux.output
)set message test on
)set message auto off
)clear all

———-

from the paper:

Does Axiom Solve Systems of O.D.E.’s Like Mathematica?

Nicolas Robidoux University of New Mexico 24 June 1993 — * —

--S 1 of 15
X1:=operator 'X1
--R
--R
--R   (1)  X1
--R                                                          Type: BasicOperator
--E 1
--S 2 of 15
deq1:=D(X1 t,t)=(1+ cos t /(2+sin t)) * X1 t
--R
--R
--R          ,     X1(t)sin(t) + X1(t)cos(t) + 2X1(t)
--R   (2)  X1 (t)= ----------------------------------
--R                            sin(t) + 2
--R                                          Type: Equation(Expression(Integer))
--E 2
--S 3 of 15
solve(deq1,X1,t)
--R
--R
--R                                 t            t
--R   (3)  [particular= 0,basis= [%e sin(t) + 2%e ]]
--RType: Union(Record(particular: Expression(Integer),basis: List(Expression(Integer))),...)
--E 3

--S 4 of 15
C1*%.basis.1
--R
--R
--R             t               t
--R   (4)  C1 %e sin(t) + 2C1 %e
--R                                                    Type: Expression(Integer)
--E 4
--S 5 of 15
function(%,'x1,'t)
--R
--R
--R   (5)  x1
--R                                                                 Type: Symbol
--E 5
--S 6 of 15
x1
--R
--R
--R   (6)  x1 t == C1 exp(t)sin(t) + 2C1 exp(t)
--R                                                     Type: FunctionCalled(x1)
--E 6
--S 7 of 15
X2:=operator 'X2
--R
--R
--R   (7)  X2
--R                                                          Type: BasicOperator
--E 7
--S 8 of 15
deq2:=D(X2 t,t)=x1 t - X2 t
--R
--R   Compiling function x1 with type Variable(t) -> Expression(Integer)
--R
--R          ,          t               t
--R   (8)  X2 (t)= C1 %e sin(t) + 2C1 %e  - X2(t)
--R
--R                                          Type: Equation(Expression(Integer))
--E 8
--S 9 of 15
solve(deq2,X2,t)
--R
--R
--R   (9)
--R                      - t   t 2                              - t   t 2

--R                2C1 %e   (%e ) sin(t) + (- C1 cos(t) + 5C1)%e   (%e )
--R   [particular= ------------------------------------------------------,
--R                                           5
--R              - t
--R    basis= [%e   ]]
--RType: Union(Record(particular: Expression(Integer),basis: List(Expression(Integer))),...)
--E 9
--S 10 of 15
%.particular
--R
--R
--R               - t   t 2                              - t   t 2
--R         2C1 %e   (%e ) sin(t) + (- C1 cos(t) + 5C1)%e   (%e )
--R   (10)  ------------------------------------------------------
--R                                    5
--R                                                    Type: Expression(Integer)
--E 10
--S 11 of 15
simplify %
--R
--R
--R               t                              t
--R         2C1 %e sin(t) + (- C1 cos(t) + 5C1)%e
--R   (11)  --------------------------------------
--R                            5
--R                                                    Type: Expression(Integer)
--E 11
--S 12 of 15
%+C2*%%(-3).basis.1
--R
--R
--R               t                              t         - t
--R         2C1 %e sin(t) + (- C1 cos(t) + 5C1)%e  + 5C2 %e
--R   (12)  --------------------------------------------------
--R                                  5
--R                                                    Type: Expression(Integer)
--E 12
--S 13 of 15
function(%,'x2,'t)
--R
--R
--R   (13)  x2
--R                                                                 Type: Symbol
--E 13
--S 14 of 15

x2
--R
--R
--R                 2C1 exp(t)sin(t) + (- C1 cos(t) + 5C1)exp(t) + 5C2 exp(- t)
--R   (14)  x2 t == -----------------------------------------------------------
--R                                              5
--R                                                     Type: FunctionCalled(x2)
--E 14

———-

We can stop here because the answer is improved over the one given in the paper. — * —

--S 15 of 15
x1 t
--R
--R
--R              t               t
--R   (15)  C1 %e sin(t) + 2C1 %e
--R                                                    Type: Expression(Integer)
--E 15
)spool
)lisp (bye)

———-

References^{[1] nothing}

7